OpenStudy (anonymous):

how would find the x-intercept of f(x)=(1/4)x^4-2x^2+1=0 how would find the x-intercept of f(x)=(1/4)x^4-2x^2+1=0 @Mathematics

OpenStudy (anonymous):

The x intercept of any equation is where y=0. So you'd want to make y (or in this case f(x)) 0 and solve.

OpenStudy (anonymous):

solve the quadratic equation \[\frac{1}{4}z^2-2z+1=0\] and once solved, replace z by \[x^2\]

OpenStudy (anonymous):

And that would be how you'd solve ^

OpenStudy (anonymous):

if it were me i would start with \[z^2-8z+4=0\] then \[z^2-8z=-4\] \[(z-4)^2=-4+16=12\] \[z-4=\pm\sqrt{12}\] \[z=4\pm2\sqrt{3}\]

OpenStudy (anonymous):

also what is the y and x intercepts for x*sqrt(2-x^(2))

OpenStudy (anonymous):

not done yet, because both these values are positive. you then have to solve \[x^2=4+2\sqrt{3}\] and \[x^2=4-2\sqrt{3}\]

OpenStudy (anonymous):

but if you want a snappier way you can write \[ (x^2-2 x-2) (x^2+2 x-2)=0\] and then solve two quadratic equation separately

OpenStudy (anonymous):

how did you get rid of the 1/4

OpenStudy (anonymous):

i can remove it if i have an equation. this is set = 0 so i can multiply by 4 and it is still = 0

OpenStudy (anonymous):

how did you get 8z+4?

OpenStudy (anonymous):

i didn't mean \[f(x) = (x^2-2 x-2) (x^2+2 x-2)\] because \[f(x) = \frac{1}{4} (x^2-2 x-2) (x^2+2 x-2)\]

OpenStudy (anonymous):

also what is the y and x intercepts for x*sqrt(2-x^(2))